# FILP-3D: Enhancing 3D Few-shot Class-incremental Learning with Pre-trained Vision-Language Models

### v1

@@ 8 Extra Ablation Studies
To further demonstrate the effectiveness of our designs, we provide more ablation results in Table 4. For no Dimensionality Reduction (DR) and no renormalized cosine similarity (RCS) case (line 1), we use cosine similarity between $\mathbf{f}^{d}$ and $\mathbf{F}^{t}$ to calculate logits. For with DR and no renormalized cosine similarity (RCS) case (line 2), we use cosine similarity between $\widetilde{\mathbf{f}}^{d}$ and $\widetilde{\mathbf{F}}^{t}$ to calculate logits. For with DR and renormalized cosine similarity (RCS) case (line 3), we use Eq.17 to calculate logits.
@@
It can be observed that if DR is used without the RCS (line 2), The improvement in performance is limited. Conversely, the combination of DR and the RCS (line 3) leads to a significant improvement in performance, particularly in terms of $\rm NCAcc$. We can thus conclude that, without using the RCS, the process of dimensionality reduction leads to inappropriate stretching, which ultimately distorts the semantic information when eliminating redundant information. This phenomenon verifies the conclusion of our analysis that redundant information affects the classification.
We can observe a significant improvement in line 3 compared to line 1 (both in $\rm Acc$ and $\rm NCAcc$). This experimental result provides evidence that redundancy within PTMs can have a pronounced impact on classification accuracy and the effectiveness of RFE in mitigating this issue.
9 Details of FSCIL3D-XL
Dataset Task Name of Classes
ShapeNet 1 airplane, ashcan, bag, basket,
bathtub, bed, bench, birdhouse,
bookshelf, bottle, bowl, bus,
cabinet, camera, can, cap,
car, cellular telephone, chair, clock,
computer keyboard, dishwasher,
display, earphone, faucet, file,
guitar, helmet, jar, knife, lamp,
laptop, loudspeaker, mailbox,
microphone, microwave,
motorcycle, mug, piano, pillow,
pistol, pot, printer, remote control,
rifle, rocket, skateboard, sofa,
stove, table, telephone, tower,
train, vessel, washer
ModelNet 2 cone, cup, curtain, desk
3 door, dresser, flower pot, glass box
4 mantel, monitor, night stand, person
5 plant, radio, range hood, sink
6 stairs, stool, tent, toilet
7 tv stand, vase, wardrobe, xbox
Table 5: Details of S2S task setup.
Dataset Task Name of Classes
ShapeNet 1 airplane, ashcan, bag, basket,
bathtub, bed, bench, birdhouse,
bookshelf, bottle, bowl, bus,
cabinet, camera, can, cap,
car, cellular telephone, chair, clock,
computer keyboard, dishwasher,
display, earphone, faucet, file,
guitar, helmet, jar, knife, lamp,
laptop, loudspeaker, mailbox,
microphone, microwave,
motorcycle, mug, piano, pillow,
pistol, pot, printer, remote control,
rifle, rocket, skateboard, sofa,
stove, table, telephone, tower,
train, vessel, washer
CO3D 2 kite, keyboard, apple, plant
3 toaster, pizza, donut, parkingmeter
4 toybus, vase, baseballglove, couch
5 broccoli, hydrant, bicycle, toilet
6 toytrain, cup, banana, sandwich
7 book, mouse, hotdog, cellphone
8 baseballbat, umbrella,
toyplane, wineglass
9 tv, orange, toytruck, ball
10 stopsign, hairdryer,
backpack, remote
11 carrot, frisbee, cake, handbag
12 suitcase
Table 6: Details of S2R task setup.
9.1 S2S Task
We choose ShapeNet [2] as our base dataset and ModelNet40 [36] as the incremental dataset. Our base task retains all 55 classes from ShapeNet [2], with 42,001 training and 10,469 test samples. For the incremental tasks, we choose 24 non-overlapping (with the base task) classes, consisting of 1,339 test instances. Table 5 shows the detailed class divisions.
9.2 S2R Task
We use ShapeNet [2] as our base dataset (similar to the S2S task, with 42,001 training and 10,469 test samples) and then choose CO3D [36] as the incremental dataset. For the incremental tasks, we exclude 9 overlapping classes that overlap with the base task, resulting in 41 distinct classes designated for the incremental tasks, consisting of 2,928 test instances. Table 6 shows the detailed class divisions.
10 Visualization and Analysis
@@
The visualization results of some classes in the test set are presented in the first six figures of Figure 4. The first three figures present the visualizations for Chowdhury et al. [3]’s model (1), ours w/o RFE (2), and ours (3) in the S2S task. The next three figures similarly display the visualizations for Chowdhury et al. [3]’s model (4), ours w/o SNC (5), and ours (6) in the S2R task. The last two figures present the visualization results of all classes in the S2R test set, with ours w/o SNC on (7) and ours on (8). These visualizations are generated using t-SNE [30].
10.1 Comparison with Chowdhury et al. [3]’s
Comparing our model with Chowdhury et al. [3]’s, three observations can be concluded: 1) In our model, the intra-class samples appear more compact, indicating that samples within the same class are closer in the feature space. In contrast, the visualizations of Chowdhury et al. [3]’s model, using chair and bed as examples, exhibit a more scattered distribution. The lack of distinct clustering centers suggests its inability to extract more generalizable features. 2) In our model, there is a notable amount of blank space, indicating that other classes have sufficient feature space for representation. In contrast, the class distributions nearly fill the entire space in the visualizations of Chowdhury et al. [3]’s model, indicating a propensity for conflicts with unrepresented classes. 3) Considering the displayed classes in Figure 4, our model can accurately distinguish between bed and bus, chair and car, while Chowdhury et al. [3]’s model demonstrates confusion as indicated by the dashed boxes.
@@ 10.2 Effect of RFE
Observing Figure 4 (2), we can notice that the bed and chair classes remain relatively dispersed compared to our model. This indicates that the model struggles to extract more generalizable features under the influence of redundant information, resulting in a more scattered distribution.
10.3 Effect of SNC
@@
The distinction between ours w/o SNC (5) and ours (6) is not apparent in specific classes. Nevertheless, by comparing ours w/o SNC (7) with ours (8) in all classes, we can observe that noise greatly affects the discriminability of features (7). The inclusion of supplementary information through SNC effectively alleviates this issue (8).
11 Details of Pre-Processing
Figure 5: Overview of pre-processing
In the pre-processing stage, we render all training samples of the base task $[\mathbf{x}^{1}_{i}]_{M}$ to generate their depth maps $[{\rm\mathbf{D}}^{1}_{i}]_{M}$ and then embed them as depth features ${\rm\mathbf{F}^{2D}}\in\mathbb{R}^{MN\times C}$, where $M$ is the number of training sample in task 1, $N$ is the number of viewpoint and $C$ is the embedding dimension. We then use the depth features ${\rm\mathbf{F}^{2D}}$ to calculate principal components $\mathbf{V}\in\mathbb{R}^{P\times C}$ through SVD, as is shown in Figure 5, where $P$ is the number of principal component.
12 Experiments on Chowdhury et al. [3]’s benchmark.
Table 7: Quantitative results on Chowdhury et al. [3]’s ShapeNet2CO3D benchmark. Bold donates the best performance, $joint$ serves solely as an upper reference limit in Chowdhury et al. [3]’s work and is not involved in the comparison.
Method Acc. in each session $\uparrow$ ${\rm\Delta}\downarrow$
0 1 2 3 4 5 6 7 8 9 10
FACT [46] 81.4 76.0 70.3 68.1 65.8 63.5 63.0 60.1 58.2 57.5 55.9 31.3
FILP-3D (ours) 85.4 81.3 80.1 77.9 75.1 73.8 72.3 70.6 69.3 67.1 65.6 23.2
Table 8: Extra quantitative results on the S2R task of FSCIL3D-XL. For each set of results, the micro/marco average are presented at the top/bottom respectively. Bold donates the best performance, $joint$ serves solely as an upper reference limit in our model and is not involved in the comparison.
Method Acc. in each session $\uparrow$ Evaluation metrics
85.4 82.0 78.2 70.4 66.9 63.3 61.9 58.1 54.8 52.1 49.2 50.3 57.0 41.1 53.4
74.2 72.1 68.8 62.3 60.1 57.9 54.5 55.1 50.9 50.8 47.5 48.7 53.2 34.4 50.9
75.1 72.6 70.4 65.4 63.3 61.0 57.3 53.9 53.6 51.4 49.6 50.3 59.7 33.0 54.6
SimpleCIL-3D (ours) 89.2 86.7 83.5 81.7 79.4 79.6 78.6 70.4 72.1 71.7 70.1 71.2 49.6 20.2 58.5
78.9 77.3 75.9 73.7 70.1 66.5 64.4 60.9 59.4 58.5 54.5 56.3 49.0 28.6 52.4
FILP-3D (ours) 90.0 87.0 86.4 85.0 83.7 82.7 81.4 79.4 78.2 76.8 74.8 74.6 60.6 17.1 66.9
79.4 75.4 75.7 72.4 70.2 68.5 65.9 63.5 62.2 59.6 57.5 57.3 59.9 27.8 58.6
As shown in Table 7, FILP-3D outperforms the other two methods. After the final task, our accuracy only drops to 65.6%, which is even close to the upper bound of Chowdhury et al. [3]’s method. However, we note that the results in this benchmark cannot distinctly indicate whether the performance degradation is caused by the new class or the old classes. We cannot conclude more observations based on it.
Unfortunately, Chowdhury et al. [3]’s work does not provide any data partitions other than ShapeNet2CO3D. Consequently, we are unable to follow them and conduct further experiments and comparisons, such as incremental within a single dataset.
13 Extra Comparison Experiments
To explain why we choose 2D PTMs as the backbone rather than selecting a 3D backbone, we compare our approach with ULIP [37] (PointMLP [14] backbone). Besides, a considerable amount of research attempts to improve CLIP’s classification ability on point clouds, such as PointCLIP [43]. We compare our method with PointCLIP in the setting of 3D FSCIL and integrate our proposed modules onto PointCLIP to validate the effectiveness and generality of our proposed modules.
Comparing ULIP with our model, it’s evident that ULIP exhibits outstanding learning capabilities in novel classes. However, the large number of trainable parameters and severely inadequate data make it extremely challenging to avoid overfitting during the FSCIL process, ultimately resulting in the highest dropping rate $\rm\Delta$ and the worst performance, except the $\rm NCAcc$ metric. Consequently, we conclude that the combination of frozen 2D PTMs pre-trained on massive data along with a simple MLP adapter is more suitable for 3D FSCIL. Conversely, ULIP, aligned only with CLIP on a small amount of data, is not as suitable for the 3D FSCIL task.
Comparing PointCLIP with our model, we observe that the inter-view adapter proposed by PointCLIP exhibits similarly inadequate performance due to its complex structure. Moreover, to make a more fair comparison, we introduce PointCLIP++ (PointCLIP backbone + RFE + SNC; the contrastive learning cannot be similarly included due to PointCLIP’s inability to specify rendering distances). Experimental results demonstrate that PointCLIP++ outperforms PointCLIP significantly across various metrics, especially in $\rm NCAcc$. This further validates the generality and effectiveness of our proposed modules.
Table 9: Quantitative results on the S2S task. For each set of results, the micro/marco average are presented at the top/bottom respectively. Bold donates the best performance, $joint$ serves solely as an upper reference limit in our model and is not involved in the comparison.
Method Pub. Year Acc. in each session $\uparrow$ Evaluation metrics
0 1 2 3 4 5 6 ${\rm NCAcc}\uparrow$ ${\rm\Delta}\downarrow$ ${\rm F}\uparrow$
FACT [46] CVPR’22 82.6 77.0 72.4 69.8 68.4 67.7 67.3 41.7 18.5 51.5
48.0 44.7 42.0 39.9 38.2 37.1 36.5 34.0 23.9 35.2
BiDist [45] CVPR’23 89.6 87.7 86.2 84.7 83.8 83.6 82.3 35.0 8.1 49.1
82.5 80.3 77.3 75.2 72.5 71.2 69.3 38.0 16.0 49.1
Chowdhury et al.’s [3] ECCV’22 86.9 84.6 82.8 78.3 78.5 71.5 68.6 50.8 21.1 58.4
73.0 62.8 65.2 62.8 60.9 57.9 55.1 45.5 24.5 49.8
SimpleCIL-3D (ours) - 90.4 88.0 85.6 85.0 84.1 78.4 80.1 68.2 11.4 73.7
78.9 76.1 72.8 71.1 70.8 70.2 69.7 68.8 11.7 69.2
FILP-3D (ours) - 90.6 89.0 86.7 84.2 83.2 81.8 82.2 79.3 9.3 80.7
80.0 76.8 74.9 72.8 71.2 70.9 70.7 77.0 11.6 73.7
$Joint$ FILP-3D - 90.6 89.1 88.5 87.8 87.4 87.6 86.8 79.9 4.2 83.2
80.0 78.5 79.0 77.2 77.5 77.7 77.3 79.9 3.4 78.6
Table 10: Quantitative results on the S2R task. For each set of results, the micro/marco average are presented at the top/bottom respectively. Bold donates the best performance, $joint$ serves solely as an upper reference limit in our model and is not involved in the comparison.
Method Acc. in each session $\uparrow$ Evaluation metrics
48.6 41.4 39.7 36.8 35.5 33.6 31.2 29.5 28.4 27.2 25.8 25.9 30.8 46.7 28.1
81.8 52.8 50.1 46.2 48.3 46.1 44.7 41.8 41.8 39.8 40.0 39.7 41.9 51.5 40.8
Chowdhury et al.’s [3] 85.2 78.6 71.0 72.0 75.2 68.8 56.1 58.5 62.9 59.1 52.2 59.4 35.3 30.3 44.3
68.2 56.2 50.5 48.4 53.5 46.7 39.9 37.6 36.9 33.1 34.3 44.1 36.5 35.3 40.0
SimpleCIL-3D (ours) 89.2 86.7 83.5 81.7 79.4 79.6 78.6 70.4 72.1 71.7 70.1 71.2 49.6 20.2 58.5
78.9 77.3 75.9 73.7 70.1 66.5 64.4 60.9 59.4 58.5 54.5 56.3 49.0 28.6 52.4
FILP-3D (ours) 90.0 87.0 86.4 85.0 83.7 82.7 81.4 79.4 78.2 76.8 74.8 74.6 60.6 17.1 66.9
79.4 75.4 75.7 72.4 70.2 68.5 65.9 63.5 62.2 59.6 57.5 57.3 59.9 27.8 58.6
$Joint$ FILP-3D 90.0 89.2 89.0 88.4 88.1 87.7 87.2 86.9 85.9 84.6 83.1 83.1 54.9 7.7 66.1
79.4 77.2 76.3 75.9 75.5 73.5 73.3 73.4 70.2 69.3 68.2 68.2 60.0 14.1 63.8
14 Effect of our proposed metrics
14.1 S2S task
By observing the experiment results presented in Table 9, we notice a substantial decline in Chowdhury et al.’s [3] performance in terms of $\rm MAcc$ when learning incrementally, coupled with poor performance in the $\rm NCAcc$ metric. This signifies their inability to effectively acquire new knowledge. Solely relying on micro-Acc. and dropping rate $\Delta$, we are unable to identify the aforementioned issues. Similarly, without $\rm NCAcc$, Bidist’s inability to effectively learn new classes also can not be revealed.
14.2 S2R task
By observing the experiment results presented in Table 10, we notice a catastrophic decline in FACT [46]’s performance in terms of $\rm MAcc$ when learning real-scanned data. This may be because The feature space trained by the manifold-mixup [31] method is relatively fragile in the aspect of structure, thus it is easy to overfit the noise information during the incremental process, especially for real-scanned data. Solely relying on metrics proposed by Chowdhury et al., we are unable to identify the aforementioned issues and then perform the analysis. Similarly, without $\rm NCAcc$, Chowdhury et al.’s terrible performance in learning real scanned new classes also can not be noticed.

### v2

@@ 6.3 Experimental Results
Table 8: Ablation studies on the S2R task of FSCIL3D-XL. The ablation experiments for the RFE module, the SNC module, and contrastive learning. were conducted in this set of experiments.
RFE SNC CL 0 11 ${\rm NCAcc}\uparrow$ ${\rm\Delta}\downarrow$ ${\rm F}\uparrow$
86.7 71.2 49.6 17.9 58.5
78.9 56.3 49.0 28.6 52.4
90.4 75.6 50.8 16.3 60.8
80.0 58.3 51.0 27.1 54.4
90.1 73.6 54.5 18.3 62.6
79.3 56.7 56.1 28.5 56.4
89.8 74.0 57.8 17.6 64.9
78.7 56.7 58.0 28.0 57.3
90.2 74.4 60.4 17.5 66.7
79.5 57.7 61.8 27.4 59.6
6.4 Ablation Studies
To evaluate the effectiveness of our proposed modules and components, we conduct several ablation experiments. We perform ablation experiments on the S2S task to investigate the contribution of the renormalized cosine similarity (RCS) and the Redundant Feature Eliminator (RFE) module. (Table 6.3) We perform ablation experiments on the S2R task to investigate the contribution of the RFE module, the spatial Noise Compensator (SNC) module, and the contrastive learning. (Table 6.3) For no contrastive learning (CL) case, training loss $\mathcal{L}^{b}=\mathcal{L}^{b}_{align}+\mathcal{L}^{b}_{cls}$
Effect of RFE and RCS. For no RFE case, we use cosine similarity between $\mathbf{f}^{g}$ and ${\rm\mathbf{F}}^{t}$ to calculate logits. In table 6.3, one can notice that RFE+SNC (line 4) significantly outperforms only SNC (line 3), this phenomenon verifies the conclusion of our analysis that redundant information affects the classification. While only RFE (line 2) can only slightly outperform baseline (line 1) in $\rm NCAcc$. We analyze the reason as follows: The huge domain gap leads to some semantic features available in synthetic data can not be extracted efficiently on real-scanned data. Adding some real-scanned data samples in the principal component extraction stage as prior knowledge may effectively alleviate this symptom.
@@ 6.4 Ablation Studies
To further demonstrate the more detailed effectiveness of RFE and RCS, we provide more ablation results in Table 6.3. For no Dimensionality Reduction (DR) and no RCS case (line 1), we use cosine similarity between $\mathbf{f}^{2D}$ and $\mathbf{F}^{t}$ to calculate logits. For with DR and no RCS case (line 2), we use cosine similarity between $\widetilde{\mathbf{f}}^{2D}$ and $\widetilde{\mathbf{F}}^{t}$ to calculate logits. For DR and renormalized cosine similarity (RCS) case (line 3), we use Eq.5 to calculate logits.
@@
In Table 6.3, it can be observed that if DR is used without the RCS (line 2), The improvement in performance is limited. Conversely, the combination of DR and the RCS (line 3) leads to a significant improvement in performance, particularly in terms of $\rm NCAcc$. We can thus conclude that, without using the RCS, the process of dimensionality reduction leads to inappropriate stretching, which ultimately distorts the semantic information when eliminating redundant information.
Effect of SNC. Let’s compare the results of the line 2 with the line 4. On the one hand, there is a slight decrease in $\rm Acc$, which indicates that the 3D module is not as capable as the multi-view model with 2D prior knowledge. It is suitable only for supplementing the information in the 3D FSCIL task. On the other hand, there is a significant increase in the incremental part of the evaluation metric, indicating that the 3D module can complement the global information that the multi-view model lacks and enhance the anti-interference ability.
Table 9: Impact of the number of views $N$.
$N$ 0 11 ${\rm NCAcc}\uparrow$ ${\rm\Delta}\downarrow$ ${\rm F}\uparrow$
4 89.6 71.0 56.4 20.8 62.9
77.6 54.5 57.5 29.8 56.0
6 90.0 74.6 60.0 17.1 66.9
79.4 57.3 59.9 27.8 58.6
8 90.0 74.1 60.0 17.7 66.3
78.4 56.7 59.3 27.7 58.0
10 90.3 75.3 62.8 16.6 68.5
79.5 57.8 63.3 27.3 60.4
14 90.0 75.0 61.8 16.7 67.8
79.4 57.1 60.7 28.1 58.8
Table 10: Impact of the batch size $BS$. Memory refers to the maximum GPU memory consumption during the experiment.
$BS$ 0 11 ${\rm NCAcc}\uparrow$ ${\rm\Delta}\downarrow$ ${\rm F}\uparrow$ Memory
8 90.0 66.1 59.6 26.5 62.7 5240Mib
78.7 56.5 59.5 28.2 58.0
16 90.0 71.8 56.7 20.2 63.4 7848Mib
79.8 57.4 58.2 28.1 57.8
32 90.0 74.6 60.0 17.1 66.9 12838Mib
79.4 57.3 59.9 27.8 58.6
64 90.2 73.8 58.9 18.2 65.5 22844Mib
79.7 57.2 60.0 28.2 58.6
128 90.3 72.8 63.7 19.6 67.9 37644Mib
79.9 56.4 64.2 29.4 60.0
Table 11: Impact of principal components $M$ of the DR
$\rm M$ of the DR 0 11 ${\rm NCAcc}\uparrow$ ${\rm\Delta}\downarrow$ ${\rm F}\uparrow$
127(80% energy) 89.6 72.1 52.4 19.5 60.7
76.9 56.9 53.3 28.6 55.0
188(90% energy) 90.2 72.4 55.7 19.7 63.0
78.6 56.1 57.9 28.6 57.0
242(95% energy) 90.0 74.6 60.0 17.1 66.9
79.4 57.3 59.9 27.8 58.6
354(99% energy) 90.5 75.0 53.7 17.1 62.6
79.9 58.4 55.5 26.9 56.9
452(99.9% energy) 90.6 75.1 54.4 17.4 63.1
80.4 58.1 55.7 27.7 56.9
Table 12: Impact of $\alpha$ and $\beta$
$\alpha$ $\beta$ 0 11 ${\rm NCAcc}\uparrow$ ${\rm\Delta}\downarrow$ ${\rm F}\uparrow$
0.1 0.1 89.9 70.5 56.3 21.6 62.6
76.9 55.6 58.4 27.7 57.0
0.1 0.25 88.8 65.2 58.1 26.6 61.4
75.8 52.0 60.1 31.4 55.8
0.25 0.1 88.9 63.2 58.7 28.9 60.9
75.0 51.1 60.3 31.9 55.3
0.25 0.25 90.2 74.4 60.4 17.5 66.7
79.5 57.7 61.8 27.4 59.7
0.25 0.5 88.8 67.2 55.1 24.3 60.6
75.4 51.4 54.5 31.1 53.0
0.5 0.25 88.6 64.9 52.6 26.7 58.1
74.7 51.5 54.5 31.1 53.0
0.5 0.5 90.0 71.8 61.1 20.2 66.0
77.6 56.4 60.1 27.3 58.2
Hyperparameter sensitivity. In Tables 6.4, 6.4, 6.4, and 6.4, we conduct experiments varying the number of views, batch size, number of principal components in dimensionality reduction (DR), and the values of $\alpha$ and $\beta$ in Eq. 12 on the S2R task. An insufficient number of views leads to a lack of semantic information, while an excessive number of views hampers the model’s ability to learn class-specific features in few-shot incremental scenarios. As a result, using an intermediate number of views, specifically 10, achieves optimal performance. However, to ensure a fair comparison, we also adopt the default setting of 6 viewpoints as used in PointCLIP [15]. A larger batch size implies improved learning capabilities, resulting in an increase in $\rm NCAcc$. However, enhanced learning capacity not only facilitates feature dispersion ($\uparrow$) but also leads to a more pronounced compression of the previous feature space ($\downarrow$), causing a nonlinear change in the model’s susceptibility to forgetting. The memory capacity of common consumer-grade GPUs, such as the NVIDIA RTX 4090 GPU, is limited to 24GB. Balancing memory cost and performance considerations, we set the batch size to 32 instead of 128. Too many principal components may fail to eliminate all redundant features, while too few principal components may remove semantic features, both of which disrupt subsequent classification. Therefore, the optimal number of principal components lies in moderation, and we set the number of principal components to 242. Regarding the selection of $\alpha$ and $\beta$, disparate learning rates between the 2D and 3D branches may cause one branch to become overfitted before the other completes alignment. If $\alpha$ and $\beta$ are too low, they may not effectively accomplish the alignment task, while excessively high values of $\alpha$ and $\beta$ may lead to the compression and excessive smoothing of 2D/3D features, resulting in a decline in model performance. Therefore, we set $\alpha$ and $\beta$ equal to 0.25.
6.5 Visualization and Analysis
@@
The visualization results of some classes in the test set are presented in figures 1-3 and 5-7 of Figure 5. Figures 1-3 present the visualizations for Chowdhury et al. [8]’s model (1), ours w/o RFE (2), and ours (3) in the S2S task. Figures 5-7 similarly display the visualizations for Chowdhury et al. [8]’s model (5), ours w/o SNC (6), and ours (7) in the S2R task. Figures 4 and 8 present the visualization results of all classes in the S2R test set, with ours w/o SNC on (4) and ours on (8). These visualizations are generated using t-SNE [41].
Comparison with Chowdhury et al. [8]’s Comparing our model with Chowdhury et al. [8]’s, three observations can be concluded: 1) In our model, the intra-class samples appear more compact, indicating that samples within the same class are closer in the feature space. In contrast, the visualizations of Chowdhury et al. [8]’s model, using chair and bed as examples, exhibit a more scattered distribution. The lack of distinct clustering centers suggests its inability to extract more generalizable features. 2) In our model, there is a notable amount of blank space, indicating that other classes have sufficient feature space for representation. In contrast, the class distributions nearly fill the entire space in the visualizations of Chowdhury et al. [8]’s model, indicating a propensity for conflicts with unrepresented classes. 3) Considering the displayed classes in Figure 5, our model can accurately distinguish between bed and bus, chair and car, while Chowdhury et al. [8]’s model demonstrates confusion as indicated by the dashed boxes.
@@
Effect of RFE Observing Figure 5 (2), we can notice that the bed and chair classes remain relatively dispersed compared to our model. This indicates that the model struggles to extract more generalizable features under the influence of redundant information, resulting in a more scattered distribution.
@@
Effect of SNC The distinction between ours w/o SNC (6) and ours (7) is not apparent in specific classes. Nevertheless, by comparing ours w/o SNC (4) with ours (8) in all classes, we can observe that noise greatly affects the discriminability of features (4). The inclusion of supplementary information through SNC effectively alleviates this issue (8).
7 Limitation
Figure 6: Most views lack informative content. Only the frontal, right-side, and top-down perspectives provide semantic information.
This work still faces limitations: first, the performance of multi-view fusion is limited when noise is prominent. For example, when point clouds are heavily occluded by walls or floors, as shown in Figure 6, less than half of the views contain semantic information. This lack of 2D information extraction and insufficient multi-view fusion further exacerbates the challenge of few-shot continual learning. Second, insufficient consideration has been given to the complementarity between 2D and 3D features, potentially leading to a suboptimal fusion strategy. For example, the integrated 2D features may mislead the 3D features when the projections of each view cannot extract effective semantic features.
8 Conclusion
In this paper, we proposed FILP-3D, a 3D few-shot class-incremental learning framework with pre-trained V-L models. Specifically, we introduced a V-L pre-trained model CLIP to the 3D FSCIL task, demonstrating its potential to mitigate overfitting and forgetting issues and narrow the domain gap in the 3D FSCIL task. To guarantee that CLIP performs well in the 3D FSCIL task, we proposed a Redundant Feature Eliminator to eliminate redundant features without stretching semantic information and a Spatial Noise Compensator to complement noise information. The two proposed modules are theoretically applicable to other 3D vision tasks utilizing V-L pre-trained models. To comprehensively evaluate models’ performance in the 3D FSCIL task, we further proposed an open-source benchmark FSCIL3D-XL, which retains all classes and introduces reasonable evaluation metrics. Extensive experiments demonstrate that our FILP-3D achieves state-of-the-art performance in all available benchmarks. Results from ablation studies and visualization further verify the effectiveness of the proposed components.
Nonetheless, this work still faces limitations, such as the multi-view fusion lacks robustness in extremely noisy scenarios, and multi-modal feature fusion is suboptimal when 2D features cannot be effectively extracted. In the future, we attempt to leverage 3D features to achieve adaptive view selection, thereby mitigating the impact of uninformative views and extracting richer 2D information. At the same time, multi-level interactive fusion and cross-modal attention can be used to explore fusion strategies, to facilitate the integration of more comprehensive global features.